The function f: R^2→R, defined by

f(x,y)= 1-x^2+y^2 has an extremum at (0,0)

True or false with full explanation


1
Expert's answer
2021-05-31T09:42:09-0400

Given f(x,y) = 1-x2+y2

differentiating f(x,y) with respect to x and y

fx(x,y) = -2x

fy(x,y) = 2y

For critical point, fx(x,y) = fy(x,y) = 0

So, x = y = 0

So critical point of f(x,y) = (0,0)

We know that partial derivative of x and y of f(x,y) is (d2f/dx2)(d2f/dy2) - (d2f/dx dy)2 = D

So, (d2f/dx2) = -2

(d2f/dy2) = 2

(d2f/dx dy) = 0

D = (-2)(2) - (0)2

D = -4 < 0

If D < 0 , then (0,0) is a saddle point.

So, (0,0) is not an extremum of f(x,y).